An auto travels at the rate of 25km/h for 4.0 minutes, then at the 50 km/h at 8.0 minutes and finally at 20 km/h at 2.0 minutes. Find (a) the total distance covered in km and (b) the average speed for the complete trip in m/s.


1
Expert's answer
2021-03-09T15:30:28-0500

(a) Let's first find the total distance traveled:


"d_{tot}=d_1+d_2+d_3,""d_{tot}=v_1t_1+v_2t_2+v_3t_3,"

"d_{tot}=25\ \dfrac{km}{h}\cdot(4\ min\cdot \dfrac{1\ h}{60\ min})+50\ \dfrac{km}{h}\cdot(8\ min\cdot \dfrac{1\ h}{60\ min})+20\ \dfrac{km}{h}\cdot(2\ min\cdot \dfrac{1\ h}{60\ min})=9\ km."

(b) We can find the average speed for the complete trip as follows:


"v_{avg}=\dfrac{d_{tot}}{t_{tot}},""v_{avg}=\dfrac{9\ km}{4\ min\cdot \dfrac{1\ h}{60\ min}+8\ min\cdot \dfrac{1\ h}{60\ min}+2\ min\cdot \dfrac{1\ h}{60\ min}},""v_{avg}=38.57\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{1\ h}{3600\ s}=10.71\ \dfrac{m}{s}."